OpenStudy (anonymous):

Simplify using exponent laws 3^(n+2) x 27^(n+1)/9^(n+1) x 3^(2n+1)

OpenStudy (zzr0ck3r):

\(\large\frac{3^{n+1}27^{n+1}}{9^{n+1}3^{2n+1}}=\frac{3^{n+1}*(3^3)^{n+1}}{(3^2)^{n+1}3^{2n+1}}=\frac{3^{n+1}3^{3n+3}}{3^{2n+2}3^{2n+1}}=\frac{3^{n+1+3n+3}}{3^{2n+2+2n+1}}\\=\frac{3^{4n+4}}{3^{4n+3}}=\frac{3^{4n}*3^4}{3^{4n}*3^3}=\frac{3^4}{3^3}=3\)

OpenStudy (anonymous):

awesome! thanks so much, it looks right

OpenStudy (zzr0ck3r):

wait

OpenStudy (zzr0ck3r):

\(\large\frac{3^{n+2}27^{n+1}}{9^{n+1}3^{2n+1}}=\frac{3^{n+2}*(3^3)^{n+1}}{(3^2)^{n+1}3^{2n+1}}=\frac{3^{n+2}3^{3n+3}}{3^{2n+2}3^{2n+1}}=\frac{3^{n+2+3n+3}}{3^{2n+2+2n+1}}\\=\frac{3^{4n+5}}{3^{4n+3}}=\frac{3^{4n}*3^5}{3^{4n}*3^3}=\frac{3^5}{3^3}=3^2=9\)

OpenStudy (zzr0ck3r):

I had 3^(n+1) we wanted 3^(n+2)

OpenStudy (anonymous):

I am a little confused on the part where you get 3^4n+5/3^4n+5.

OpenStudy (anonymous):

How did you get 5? or rather 2+3 before that

OpenStudy (anonymous):

Do you multiply (3^3) to both n and +1?